OpenStudy (anonymous):

The depth (inches) of water that accumulates in the spring soil from melted snow can be determined by dividing the cumulative winter snowfall (inches) by 12. Determine the amount of winter snow that accumulates 6 inches of water in the soil.

OpenStudy (anonymous):

12∗6=72/12=6 this should help im not exactly positive though

OpenStudy (anonymous):

12*6=72/12=6

OpenStudy (anonymous):

?? did i help??

OpenStudy (anonymous):

yupp